OpenStudy (anonymous):

Solve -3 (6x - 4) + 11x = -9 Please dont explain to me just do it and show your work i will fan you and give you a medal Because I only have a few more minuts to turn i in please do alll the steps thanks

OpenStudy (anonymous):

@subway_surfers

OpenStudy (anonymous):

are you solving for x

OpenStudy (anonymous):

i guys i am not reallyt very sure

OpenStudy (anonymous):

x= 3

OpenStudy (anonymous):

i know bu i need the work shown

OpenStudy (anonymous):

-3 (6x - 4) + 11x = -9 distribute -3 (6x - 4) it becomes -18x + 12 + 11x = -9. combine like terms it becomes -7x + 12 = -9 now its a simple two step equation. subtract 12 from both sides to get -7x = -21 divide by -7 on both sides to get your answer : x = 3

OpenStudy (anonymous):

-18x+12+11x+=-9 -7x=-21 x=-3

OpenStudy (anonymous):

timotheos you're wrong its positive 3

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

hey just in 2nd grade

OpenStudy (anonymous):

loll ya right

OpenStudy (anonymous):

but sam do you get it

OpenStudy (anonymous):

really you want to see my photo

OpenStudy (anonymous):

um. no

OpenStudy (anonymous):

awkward.......

OpenStudy (anonymous):

yay thx

OpenStudy (anonymous):

gtg doing Logarithms

OpenStudy (anonymous):

um sam wtf LOL